OpenStudy (anonymous):

-3/5*[2/3-1/2(3/4-1/2)]=

OpenStudy (anonymous):

- 3/5

OpenStudy (anonymous):

\[-\frac{3}{5} \left(\frac{2}{3}-\frac{1}{2} \left(\frac{3}{4}-\frac{1}{2}\right)\right) = -\frac{13}{40} \]